OpenStudy (anonymous):

The current in a certain circuit as measured by an ammeter is a continuous random variable X with the following density function:

OpenStudy (anonymous):

f(x) = {0.075x + 0.2 3≤x≤5 { 0 otherwise Calculate P(X≤4). How does this probability compare to P(X<4)? I calculated (0.075x + 0.2)dx on the interval of 3 to 4. The answer in the back of the book was 0.25, but is this really correct?

OpenStudy (loser66):

I don't know. To me, \(P(x\leq 4)\) is the same as \(P(x<4)\) And my answer is 0.4625, not 0.25 :(

OpenStudy (loser66):

@Zarkon

OpenStudy (anonymous):

Yeah, that's the same answer I got. Hrm...

OpenStudy (kropot72):

My result is the same:\[\large P(X \le4)=0.4625\]

OpenStudy (zarkon):

For any continuous rv \(X\) we have \(P(X=a)=0\) for any constant \(a\) therefore \[P(X\le 4)=P(X<4 \text{ or } X=4)\] \[=P(X<4)+P(X=4)=P(X<4)+0=P(X<4)\]

OpenStudy (loser66):

Yes, now it is clear for why \(P(x\leq 4)\) is same as \(P(x< 4)\). How about the answer? How can they get 0.25?

OpenStudy (zarkon):

it is a mistake

OpenStudy (zarkon):

either RightInTheCranium looked at the wrong answer or the answer in the book is not correct (or the problem was typed wrong)

OpenStudy (loser66):

Yeaaaaaaaah!! If it is not, I do not want to study statistics any more. :)

OpenStudy (loser66):

Thanks a lot @Zarkon

OpenStudy (anonymous):

Double and triple checked. Back of the book says 0.5. Thanks for the help though! Glad to know I'm not just spinning my wheels.